OpenStudy (anonymous):

the perrimeter of a standard- sized rectangular rug is 24 ft. The length is 2 ft longer than the width. find the dimensions

myininaya (myininaya):

\[2L+2W=24\] \[L=2+W\]

myininaya (myininaya):

we shall plug equation 2 into 1 so we have \[2(2+W)+2W=24 => 4+2W+2W=24=>4+4W=24 => 4W=20 =>W=5\]

myininaya (myininaya):

\[L=2+5=7\]

OpenStudy (anonymous):

This is a simple substitution problem. 2L + 2W == 24 As L == W + 2, rewrite this as 2W + 2(W + 2) == 24 Then, 2W + 2W + 4 == 24 4W + 4 == 24

OpenStudy (anonymous):

let width=x ft then length=2+x hence perimeter of rectangle - 2(length+width)=24 =>2(2+x+x)=24 =>2x=10 =>x=5 ft hence length=7ft and width=5 ft